OpenStudy (yanni):

In an electrical circuit, the current passing through a conductor varies inversely with the resistance. Suppose that when the current is 10A(amperes), the resistance is 14 ohms. What is the current when the resistance is 5 ohms?

OpenStudy (shubhamsrg):

well i = k/r (k=constant) thus ir = constant so i1 * r1 = i2 * r2 .. hope this'll help

OpenStudy (stormfire1):

Using standard ohm's-law speak. I = current, R = resistance, V = voltage: \[10A * 14 \Omega = 140V\]\[ I * 5 \Omega= 140V\]\[I=28A\]